OpenStudy (anonymous):

x^2 - 2x - 35? fatorise?

jimthompson5910 (jim_thompson5910):

x^2 - 2x - 35 x^2 - 7x+5x - 35 (x^2 - 7x)+(5x - 35) x(x-7)+5(x-7) (x+5)(x-7) So x^2 - 2x - 35 completely factors to (x+5)(x-7)

OpenStudy (jamesj):

And the way Jim figured that out was "what are the factors of -35 that add up to 2?

OpenStudy (jamesj):

and that's because if a quadratic has factors x-a and x-b, then (x-a)(x-b) = x^2 - (a+b)x + ab. Hence with your equation x^2 - 2x - 35 you know that a+b = 2; and ab = -35. Hence we want factors of -35 that add up to 2. Those factors are +7 and -5.